# Scene-Level Appearance Transfer with Semantic Correspondences

### v1

@@ 3.2 Appearance Transfer via Semantic Matching
We leverage the open-vocabulary panoptic segmentation model ODISE [63] for semantic matching. For a given input image, ODISE generates segmentation maps $\mathcal{M}\in\{1,\ldots,C\}^{H\times W}$, assigning each pixel to one of $C$ semantic categories. These maps enable semantic correspondences between the style and source images (detailed below). By matching open-vocabulary semantic predictions, ReStyle3D is not limited by predefined semantic categories in a scene. The correspondences are injected into the diffusion process to guide appearance transfer between matched regions.
Injecting Semantic Correspondences in Self-attention.
ReStyle3D enables training-free style transfer by extending the self-attention layer of a pretrained diffusion model (Fig. 2). This approach injects style information from $\mathbf{I}_{style}$ into $\mathbf{I}_{src}$ while preserving its structure. Specifically, we first encode both the style and source images into the latent space of Stable Diffusion [49], producing $\mathbf{z}^{style}_{0}$ and $\mathbf{z}^{src}_{0}$. These latent representations are then inverted to Gaussian noise, $\mathbf{z}^{style}_{T}$ and $\mathbf{z}^{src}_{T}$, using edit-friendly DDPM inversion [25]. To enhance structural preservation and mitigate LDM’s over-saturation artifacts, we incorporate monocular depth estimates [64] of the input images through a depth-conditioned ControlNet [71] during the inversion process. The stylized image latent is then initialized as $\mathbf{z}^{out}_{T}=\mathbf{z}^{src}_{T}$.
Next, we transfer the style from $\mathbf{z}^{style}_{T}$ to $\mathbf{z}^{out}_{T}$ by de-noising them along parallel paths [1]. At each de-noising step $t$, we extract style features $(K_{style},V_{style})$ and query features $Q_{out}$ from individual self-attention layers. The semantic-guided attention for the output feature $\phi_{out}$ is computed by combining the attention features with the attention mask $M$ as follows:
$\phi_{out}=\text{softmax}\left(\frac{Q_{out}\cdot K_{style}^{T}}{\sqrt{d_{h}}}\odot M\right)\cdot V_{style},$ (4)
@@
To obtain the attention mask $M\in\mathbb{R}^{d^{2}\times d^{2}}$, we flatten and bilinearly downsample the semantic masks $\mathcal{M}_{style}$ and $\mathcal{M}_{src}$ to match the resolution of attention feature maps, which is $d\times d$. The attention mask is defined as $M(i,j)=1$ if the $i$-th region in the source and the $j$-th region in the style image share the same semantic class; otherwise, $M(i,j)=0$. This formulation ensures that each region in the output image samples its appearance solely from semantically corresponding regions in the style image. For example, a chair in the source image is only cross-attended to its counterpart in the style image, inheriting its appearance. If multiple instances in the style image share the same semantic class, attention is distributed across them based on sampling weights determined by softmax attention scores. This mechanism naturally extends to support user-specified correspondences. Regions without semantic matches attend to the entire style image to preserve global harmony. Although semantic attention effectively transfers appearance, it may compromise the realism and structure of the stylized output, requiring further refinement.
@@
We draw inspiration from [22, 1] and incorporate classifier-free guidance (CFG) combined with semantic and depth-conditioned generation. At each denoising step $t$, we compute three noise predictions: $\epsilon_{t}$, $\epsilon_{t}^{d}$ and $\epsilon_{t}^{s}$. Here, $\epsilon_{t}^{s}$ represents the predicted noise from the semantic attention path, $\epsilon_{t}^{d}$ is obtained from the depth-conditioned ControlNet [71], and $\epsilon_{t}$ is the unconditional noise prediction. The final noise prediction is then calculated as follows:
$\hat{\epsilon}_{t}=(1-\alpha)\epsilon_{t}+\alpha(\lambda_{s}\epsilon_{t}^{s}+\lambda_{d}\epsilon_{t}^{d}),$ (5)
@@
To enhance image quality, we employ a two-stage refinement process. First, we upscale the initial stylized image from 512×512 to 1024×1024 resolution. Then, following SDEdit [38], we add high-frequency noise to this upscaled image and denoise it for 100 steps with SDXL [46]. This refinement process enhances local details while maintaining the overall style, producing our final output $\hat{\mathbf{I}}_{src}$.
Figure 3: Appearance Transfer Multi-view Inconsistency. When stylizing each view separately, we observe inconsistencies (red arrows) due to high variance in generative modeling.
Figure 4: Multi-view Style Lifting. Stereo correspondences are extracted from the original image pair $(\mathbf{I}_{src}^{i},\mathbf{I}_{src}^{j})$ and used to warp the stylized image $\hat{\mathbf{I}}^{i}$ to the second image, $\mathbf{I}^{j}_{w}$. To address missing pixels from warping, we train a warp-and-refine model to complete the stylized image $\hat{\mathbf{I}}^{j}$. This model is applied across multiple views within our auto-regressive framework.
3.3 Multi-view Consistent Appearance Transfer
Although our semantic attention module effectively transfers appearance for a single view, independently applying it to each view may cause inconsistent artifacts (see Fig. 3). Therefore, we develop an approach to transfer the appearance from the stylized image $\hat{\mathbf{I}}_{src}^{i}$ to all remaining views while maintaining multi-view consistency.
@@ 3.3 Multi-view Consistent Appearance Transfer
Given a pair of source images $(\mathbf{I}_{src}^{i},\mathbf{I}_{src}^{j})$, we first leverage a stereo matching method DUSt3R [60] to extract the dense point correspondence and the camera intrinsics. Using these, the optical flow $\mathbf{W}_{i\rightarrow j}\in\mathbb{R}^{H\times W\times 2}$ is calculated by projecting the pointmaps of $i$-th image to the $j$-th image. Next, given the optical flow and the stylized $i$-th image $\hat{\mathbf{I}}_{src}^{i}$, we employ softmax splatting [41] to obtain the initial stylized image $\hat{\mathbf{I}}_{w}^{j}$ and its warping mask $\mathbf{M}_{w}^{j}$, which indicates missing pixels in the $j$-th frame after forward warping.
Table 1: Quantitative comparison of ReStyle3D and baseline methods on 2D appearance transfer. Our method achieves the best overall performance for both structure preservation and percetual similarity, benefiting from its explicit semantic guidance and refinement.
@@
Cross-Image-Attn. [1] 22.47 7.944 5.78 0.553 0.709 0.414 4.8
IP-Adatper SDXL [66] 9.38 1.847 79.29 0.570 0.752 0.371 3.0
StyleID [7] 11.25 2.59 93.44 0.546 0.741 0.332 3.2
@@
Given the source image $\mathbf{I}_{src}^{j}$ and its initial stylized version $\hat{\mathbf{I}}_{w}^{j}$, we train a 2-view warp-and-refine model $\hat{\epsilon}_{\theta}=\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,c)$ to generate a complete and consistent stylized image following conditions $c$: the initial stylized image, the inpainting mask, and the monocular depth map $\mathbf{D}^{j}$ of the source image $\mathbf{I}_{src}^{j}$ (Fig. 4). The final condition $c=\mathrm{concat}(\mathbf{z}^{(\hat{\mathbf{I}}_{w}^{j})},\mathbf{z}^{(\mathbf{M}_{w}^{j})},\mathbf{z}^{(\mathbf{D}^{j})})$, $\mathbf{z}^{*}$ denotes individual latent representations. To harness the power of a pretrained diffusion model [46], like Marigold [27], we modify the input channels of its initial convolution layer to accommodate additional conditions and zero-initializing the additional weights. Following Eq. (2), we train the model using quadruplets of the warped and incomplete image, depth map, mask, and the clean and complete image. The model simultaneously learns to complete missing pixels and refine all pixels to address warping artifacts.
@@
We propose an auto-regressive approach to extend two-view stylization to handle multiple views or even videos, ensuring global coherence across the scene (Fig. 4). Stylizing the $j$-th frame using only the previous frame $(j-1)$ can lead to inconsistencies with earlier frames while warping all historical frames could produce blurry outputs. Instead, we warp the stylized frame $(j-1)$ along with two randomly selected historical frames. In overlapping regions, where multiple pixels are warped to the same location, we adopt an exponential weighted averaging to blend pixels, prioritizing pixels from frame $(j-1)$. This adaptive weighting maintains temporal consistency and preserves sharp details in the resulting warped image $\hat{\mathbf{I}}_{w}^{1:j-1}$. Finally, our model refines the output, producing a fully stylized frame.
4 Experiments
@@
Figure 5: Image appearance transfer results. Our method enables precise appearance transfer between semantically corresponding elements, evidenced by the green rug and glass table (first row), textured cabinet (second row), and bedsheets (third row). Unlike baselines that either apply global style transfer or fail to preserve structure, ReStyle3D maintains both semantic fidelity and structural integrity.
@@
We base our semantic attention module on Stable Diffusion 1.5 [49] and the refinement and 2-view warp-and-refine model on SDXL [46]. To train our two-view warp-and-refine model (Sec. 3.3), we use 4 NVIDIA A100 40GB GPUs with an effective batch size of 256 for 20K iterations, using the AdamW optimizer [36] with learning rate $10^{-4}$. We randomly drop out half of the text prompt during training to make our model agnostic to text conditions. The model is trained on a dataset with 57K house tour images featuring 57 different houses/apartments.
4.1 Evaluation Setting
@@
We evaluate multiple different aspects of our pipeline. First, we assess the appearance transfer performance using source images on two aspects: structure preservation and style transfer quality. For structure preservation, we compare depth maps predicted by DepthAnythingV2 [64] between stylized and original images using standard metrics: Absolute Relative Error (AbsRel), $\delta 1$ accuracy, and Squared Relative Error (SqRel), following established protocols [27, 64]. For style transfer quality, we measure perceptual similarity between the stylized output and the style image using DINOv2 [42], CLIP, and DreamSim [15] scores. We evaluate this task on the stylized source images of each scene. Next, we evaluate our two-view lifting model (Sec. 3.3). We assess its warp-and-refine quality using PSNR, SSIM [61], and LPIPS [72] while also reporting FID [21] to quantify the realism of generated frames under challenging viewpoint extrapolation. We evaluate using pairs of the source images per scene and their warped projections on the rest of the frames in each scene—we exclude pairs without correspondences. We do not use any stylization to train or evaluate since there is no ground truth. To evaluate global consistency, we leverage DUSt3R [60] to extract poses by aligning point maps from stylized sequences and compute cumulative error curve (AUC) by comparing recovered camera poses against those from original images.
4.2 Results
@@
We compare with three state-of-the-art methods on image-conditioned stylization and appearance transfer: Cross Image Attention [1], IP-Adapter [66], and StyleID [7]. For a fair comparison, we add depth ControlNet [71] to SDXL IP-Adapter [66] and use the style image as the image prompt. As shown in Tab. 1, our method achieves superior performance on both structure preservation and style transfer metrics. Notably, our explicit semantic attention mechanism in the diffusion UNet enhances the perceptual similarity between stylized outputs and style images, as evidenced by better DINO, CLIP, and DreamSim scores. The refinement step further improves structure preservation, reducing AbsRel from 11.30 to 8.34 and SqRel from 2.65 to 1.67. Qualitative comparisons (Figs. 5 and 7) reveal the limitations of existing approaches. Cross Image Attention effectively captures style textures but fails to maintain scene structure due to the lack of semantic guidance. IP-Adapter SDXL preserves overall structure but struggles with local detail transfer, as it compresses style information into a global feature vector. Although StyleID achieves the second-best performance, its results tend to preserve high-frequency details from the source image while applying style changes more globally, demonstrating limited capability in fine-grained appearance transfer.
We conduct a user study with 27 participants who were shown examples of a source and style image with outputs from four methods. Participants selected the result that best preserved the structure while faithfully transferring the style. Out of 252 evaluations (Tab. 2), ReStyle3D was the most preferred (42.4%), demonstrating its effectiveness in balancing structure preservation and appearance transfer under human perception.
Table 2: Image Appearance Transfer User Study. We show user preference rates (%) for different methods, where participants selected the result that best preserved the original scene structure while closely matching the reference style. ReStyle3D achieves the highest preference rate.
@@
Table 3: Results on two-view novel-view synthesis. ReStyle3D achieves the highest scores on all metrics, indicating more accurate view synthesis and visually pleasing outputs compared to existing methods.
@@
GenWarp [50] 5122 13.503 0.465 0.435 59.965
SDXL Inpainting [46] 16.228 0.535 0.389 89.502
ViewCrafter [69] 17.178 0.594 0.278 56.127
@@
GenWarp [50] 10242 13.491 0.565 0.440 60.540
SDXL Inpainting [46] 16.153 0.565 0.426 89.537
ViewCrafter [69] 17.137 0.652 0.317 57.898
@@
Figure 6: Results on Video/Multi-view Appearance Transfer of ReStyle3D. We show the style images, three frames stylized by ReStyle3D, followed by a 3D reconstruction of these outputs using an off-the-shelf pipeline. Despite challenging camera motion and multiple objects in the scene, our method preserves consistent geometry and seamlessly transfers the reference style across all frames.
@@
We compare our approach to: i) SDXL inpainting model [46] with depth-conditioned ControlNet [71], ii) GenWarp [50], an image-based diffusion model for single view NVS, and iii) ViewCrafter [69], a video-diffusion model for NVS. Note that the proposed task differs from traditional NVS as it leverages geometry information from the novel view itself. We employ DUSt3R [60] to extract the correspondences and provide the initial warped image as input to all methods. ReStyle3D outperforms across all metrics, achieving a superior reconstruction ability as evidenced by the best PSNR, SSIM, and LPIPS metrics (cf. Tab. 3, ). Additionally, it exhibits strong capability in extending style to unseen regions, evidenced by the lowest FID score (Fig. 8). Notably, the second best method ViewCrafter [69], requires a predefined camera trajectory as input to video diffusion and runs 10$\times$ slower than ours.
Table 4: Pose deviation from real-world estimates. We measure the fraction of camera poses within certain rotation (at $5^{\circ}$, $10^{\circ}$, $15^{\circ}$) and translation (at $1,\mathrm{cm}$, $2,\mathrm{cm}$, $5,\mathrm{cm}$) error thresholds, reporting area-under-curve (AUC) values. ReStyle3D achieves significantly higher AUC in both, showing superior multi-view geometric consistency vs. existing methods.
@@
GenWarp [50] 25.89 46.70 58.89 58.38 59.39 70.05
SDXL Inpainting [46] 34.52 52.79 66.50 61.42 65.99 74.11
ViewCrafter [69] 37.56 55.33 68.53 60.91 65.99 77.16
@@
We further evaluate the multi-view consistency of the stylized results through a proxy task. Specifically, we input the original and stylized images to DUSt3R [60] and estimate the camera poses, separately. By evaluating the agreement with the poses from the original images, we analyze whether the geometry is preserved in the stylized images. The results are presented in Tab. 4. Enabled by our adaptive auto-regressive approach, which effectively mitigates inconsistencies while preserving image sharpness, our method significantly outperforms the baselines on both rotation and translation metrics. Figs. 8 and 9 show multi-view transfer results, including the 3D reconstruction of stylized outputs with estimated camera poses, demonstrating both geometric and stylistic consistency despite camera motion and multiple objects.
@@
In Tab. 5(a), we run ReStyle3D without our guidance strategy and observe significant degradation in structure preservation (AbsRel↑ from 8.34 to 16.72). In (b), removing semantic attention hurts performance on perceptual similarity w.r.t. style image, showing that both components are crucial for semantic-accurate style transfer while maintaining structural integrity.
Table 5: Ablation Study. We separately remove the guidance strategy and the semantic attention module to evaluate their impact on both structure preservation and style fidelity. Removing either significantly degrades performance, highlighting the importance of both components in achieving robust scene geometry and perceptually faithful stylization.

### v2

@@ 3.2. Appearance Transfer via Semantic Matching
We leverage the open vocabulary panoptic segmentation model ODISE (Xu et al., 2023) for semantic matching. For a given input image, ODISE generates segmentation maps $\mathcal{M}\in\{1,\ldots,C\}^{H\times W}$, assigning each pixel to one of $C$ semantic categories. These maps enable semantic correspondences between the style and source images (detailed below). By matching open-vocabulary semantic predictions, ReStyle3D is not limited by predefined semantic categories in a scene. The correspondences are injected into the diffusion process to guide appearance transfer between matched regions.
Injecting Correspondences in Self-attention.
ReStyle3D enables training-free style transfer by extending the self-attention layer of a pretrained diffusion model (Fig. 2). This approach injects style information from $\mathbf{I}_{style}$ into $\mathbf{I}_{src}$ while preserving its structure. Specifically, we first encode both the style and source images into the latent space of Stable Diffusion (Rombach et al., 2022), producing $\mathbf{z}^{style}_{0}$ and $\mathbf{z}^{src}_{0}$. These latent representations are then inverted to Gaussian noise, $\mathbf{z}^{style}_{T}$ and $\mathbf{z}^{src}_{T}$, using edit-friendly DDPM inversion (Huberman-Spiegelglas et al., 2024). To enhance structural preservation and mitigate LDM’s over-saturation artifacts, we incorporate monocular depth estimates (Yang et al., 2024) of the input images through a depth-conditioned ControlNet (Zhang et al., 2023c) during the inversion process. The stylized image latent is then initialized as $\mathbf{z}^{out}_{T}=\mathbf{z}^{src}_{T}$.
Next, we transfer the style from $\mathbf{z}^{style}_{T}$ to $\mathbf{z}^{out}_{T}$ by de-noising them along parallel paths (Alaluf et al., 2024). At each de-noising step $t$, we extract style features $(K_{style},V_{style})$ and query features $Q_{out}$ from individual self-attention layers. The semantic-guided attention for the output feature $\phi_{out}$ is computed by combining the attention features with the attention mask $M$ as follows:
(4) $\phi_{out}=\text{softmax}\left(\frac{Q_{out}\cdot K_{style}^{T}}{\sqrt{d_{h}}}\odot M\right)\cdot V_{style},$
@@
To obtain the attention mask $M\in\mathbb{R}^{d^{2}\times d^{2}}$, we flatten and bilinearly downsample the semantic masks $\mathcal{M}_{style}$ and $\mathcal{M}_{src}$ to match the resolution of attention feature maps, which is $d\times d$. The attention mask is defined as $M(i,j)=1$ if the $i$-th region in the source and the $j$-th region in the style image share the same semantic class; otherwise, $M(i,j)=0$. This formulation ensures that each region in the output image samples its appearance solely from semantically corresponding regions in the style image. For example (Figure 3), a rug in the source image is only cross-attended to its counterpart in the style image, inheriting its appearance. If multiple instances in the style image share the same semantic class, attention is distributed across them based on sampling weights determined by softmax attention scores. This mechanism naturally extends to support user-specified correspondences. Regions without semantic matches attend to the entire style image to preserve global harmony. While semantic attention effectively transfers appearance, it may compromise realism and structure, requiring further refinement.
Figure 3. Attention Query Visualization. We visualize the attention score at two query positions, coffee table and rug. Raw attention in (Alaluf et al., 2024) spilled across regions (red arrows) due to multi-instance ambiguity, semantic attention effectively confines the activation in the matched region.
@@
We draw inspiration from (Ho and Salimans, 2021; Alaluf et al., 2024) and incorporate classifier-free guidance (CFG) combined with semantic and depth-conditioned generation. At each denoising step $t$, we compute three noise predictions: $\epsilon_{t}$, $\epsilon_{t}^{d}$ and $\epsilon_{t}^{s}$. Here, $\epsilon_{t}^{s}$ represents the predicted noise from the semantic attention path, $\epsilon_{t}^{d}$ is obtained from the depth-conditioned ControlNet (Zhang et al., 2023c), and $\epsilon_{t}$ is the unconditional noise prediction. The final noise prediction is then calculated as follows:
(5) $\hat{\epsilon}_{t}=(1-\alpha)\epsilon_{t}+\alpha(\lambda_{s}\epsilon_{t}^{s}+\lambda_{d}\epsilon_{t}^{d}),$
@@
To enhance image quality, we employ a two-stage refinement process. First, we upscale the initial stylized image from 512×512 to 1024×1024 resolution. Then, following SDEdit (Meng et al., 2022), we add high-frequency noise to this upscaled image and denoise it for 100 steps with SDXL (Podell et al., 2024). This refinement process enhances local details while maintaining the overall style, producing our final single-view output $\hat{\mathbf{I}}_{src}$.
Figure 4. Multi-view Inconsistency Caused by Separate Transfer. When stylizing each view separately, we observe inconsistencies in the results (highlighted by red arrows) due to high variance in generative modeling.
Figure 5. Multi-view Style Lifting. Stereo correspondences are extracted from the original image pair $(\mathbf{I}_{src}^{i},\mathbf{I}_{src}^{j})$ and used to warp the stylized image $\hat{\mathbf{I}}^{i}$ to the second image, $\mathbf{I}^{j}_{w}$. To address missing pixels from warping, we train a warp-and-refine model to complete the stylized image $\hat{\mathbf{I}}^{j}$. This model is applied across multiple views within our auto-regressive framework.
3.3. Multi-view Consistent Appearance Transfer
While our semantic attention module effectively transfers the appearance for a single view, applying this independently to each view may cause inconsistent artifacts (see Fig 4). Thereby, we develop an approach to transfer the appearance of the stylized image $\hat{\mathbf{I}}_{src}^{i}$ to all remaining views, while maintaining multi-view consistency.
@@ 3.3. Multi-view Consistent Appearance Transfer
Given a pair of source images $(\mathbf{I}_{src}^{i},\mathbf{I}_{src}^{j})$, we first leverage the stereo matching method (Wang et al., 2024) to extract the dense point correspondence and the camera intrinsics. Using these, the optical flow $\mathbf{W}_{i\rightarrow j}\in\mathbb{R}^{H\times W\times 2}$ is calculated by projecting the pointmaps of $i$-th image to the $j$-th image. Next, given the optical flow and the stylized $i$-th image $\hat{\mathbf{I}}_{src}^{i}$, we employ softmax splatting (Niklaus and Liu, 2020) to obtain the initial stylized image $\hat{\mathbf{I}}_{w}^{j}$ and its warping mask $\mathbf{M}_{w}^{j}$, which indicates missing pixels in the $j$-th frame after forward warping.
Table 1. Quantitative Comparison of ReStyle3D and Baseline Methods on 2D Appearance Transfer. Our method achieves the best overall performance for both structure preservation and perceptual similarity, benefiting from its explicit semantic guidance and two-stage refinement.
@@
Cross-Image-Attn. (Alaluf et al., 2024) 22.47 7.944 5.78 0.553 0.709 0.414 4.8
IP-Adatper SDXL (Ye et al., 2023) 9.38 1.847 79.29 0.570 0.752 0.371 3.0
StyleID (Chung et al., 2024) 11.25 2.59 93.44 0.546 0.741 0.332 3.2
@@
Given the source image $\mathbf{I}_{src}^{j}$ and its initial stylized version $\hat{\mathbf{I}}_{w}^{j}$, we train a 2-view warp-and-refine model $\hat{\epsilon}_{\theta}=\hat{\epsilon}_{\theta}(\mathbf{z}_{t},t,c)$ to generate a complete and consistent stylized image following conditions $c$: the initial stylized image, the inpainting mask, and the monocular depth map $\mathbf{D}^{j}$ of the source image $\mathbf{I}_{src}^{j}$ (Fig. 5). The final condition $c=\mathrm{concat}(\mathbf{z}^{(\hat{\mathbf{I}}_{w}^{j})},\mathbf{z}^{(\mathbf{M}_{w}^{j})},\mathbf{z}^{(\mathbf{D}^{j})})$, $\mathbf{z}^{*}$ denotes individual latent representations. To harness the power of a pretrained diffusion model (Podell et al., 2024), like (Ke et al., 2024), we modify the input channels of its initial convolution layer to accommodate additional conditions and zero-initializing the additional weights. Following Eq. (2), we train the model using quadruplets of the warped and incomplete image, depth map, mask, and the clean and complete image. The model learns to complete missing pixels, while globally refining all pixels to address warping artifacts.
@@
We propose an autoregressive approach to extend two-view stylization to handle multiple views or even videos, ensuring global coherence across the scene (Fig. 5). Stylizing the $j$-th frame using only the previous frame $(j-1)$ can lead to inconsistencies with earlier frames while warping all historical frames could produce blurry outputs. Instead, we warp the stylized frame $(j-1)$ along with two randomly selected historical frames. In overlapping regions, where multiple pixels are warped to the same location, we adopt an exponential weighted averaging to blend pixels, prioritizing pixels from frame $(j-1)$. This adaptive weighting ensures temporal consistency while preserving sharp details in the resulting warped image $\hat{\mathbf{I}}_{w}^{1:j-1}$. Finally, our model refines the output, producing a fully stylized frame.
4. Experiments
@@
Figure 6. Image Appearance Transfer Results. Our method enables precise appearance transfer between semantically corresponding elements, evidenced by the green rug and glass table (first row), textured cabinet (second row), and bedsheets (third row). Unlike baselines that either apply global style transfer or fail to preserve structure, ReStyle3D maintains both semantic fidelity and structural integrity.
@@
We base our semantic attention module on Stable Diffusion 1.5 (Rombach et al., 2022) and the refinement and 2-view warp-and-refine model on SDXL (Podell et al., 2024). To train our two-view warp-and-refine model (Sec. 3.3), we use 4 NVIDIA A100 40GB GPUs with an effective batch size of 256 for 20K iterations, using the AdamW optimizer (Loshchilov and Hutter, 2019) with learning rate $10^{-4}$. We randomly drop out half of the text prompt during training to make our model agnostic to text conditions. The model is trained on a dataset with 57K house tour images featuring 57 different houses and apartments.
4.1. Evaluation Setting
@@
We evaluate multiple different aspects of our pipeline. First, we assess the appearance transfer performance using source images on two aspects: structure preservation and style transfer quality. For structure preservation, we compare depth maps predicted by DepthAnythingV2 (Yang et al., 2024) between stylized and original images using standard metrics: Absolute Relative Error (AbsRel), $\delta 1$ accuracy, and Squared Relative Error (SqRel), following established protocols (Ke et al., 2024; Yang et al., 2024). For style transfer quality, we measure perceptual similarity between the stylized output and the style image using DINOv2 (Oquab et al., 2024), CLIP, and DreamSim (Fu et al., 2023) scores. We evaluate this task on the stylized source images of each scene. Next, we evaluate our two-view lifting model (Sec. 3.3). We assess its warp-and-refine quality using PSNR, SSIM (Wang et al., 2004), and LPIPS (Zhang et al., 2018) while also reporting FID (Heusel et al., 2017) to quantify the realism of generated frames under challenging viewpoint extrapolation. We evaluate using pairs of the source images per scene and their warped projections on the rest of the frames in each scene—we exclude pairs without correspondences. We do not use any stylization to train or evaluate since there is no ground truth. To evaluate global consistency, we leverage DUSt3R (Wang et al., 2024) to extract poses by aligning point maps from stylized sequences and compute cumulative error curve (AUC) by comparing recovered camera poses against those from original images.
4.2. Results
@@
We compare with three state-of-the-art methods on image-conditioned stylization and appearance transfer: Cross Image Attention (Alaluf et al., 2024), IP-Adapter (Ye et al., 2023), and StyleID (Chung et al., 2024). For a fair comparison, we add depth ControlNet (Zhang et al., 2023c) to SDXL IP-Adapter (Ye et al., 2023) and use the style image as the image prompt. As shown in Tab. 1, our method achieves superior performance on both structure preservation and style transfer metrics. Notably, our explicit semantic attention mechanism in the diffusion UNet enhances the perceptual similarity between stylized outputs and style images, as evidenced by better DINO, CLIP, DreamSim scores, and attention visualization (Fig. 3). The refinement step further improves structure preservation, reducing AbsRel from 11.30 to 8.34 and SqRel from 2.65 to 1.67. Qualitative comparisons (Figs. 6 and 9) reveal the limitations of existing approaches. Cross Image Attention effectively captures style textures but fails to maintain scene structure due to the lack of semantic guidance. IP-Adapter SDXL preserves overall structure but struggles with local detail transfer, as it compresses style information into a global feature vector. While StyleID achieves the second-best performance, its results tend to keep high-frequency details from the source image while applying style changes more globally, showing limited capability in fine-grained appearance transfer.
We conduct a user study with 27 participants who were shown examples of a source and style image with outputs from four methods. Participants selected the result that best preserved the structure while faithfully transferring the style. Out of 252 evaluations (Tab. 2), ReStyle3D was the most preferred, achieving the highest preference (42.4%) and demonstrating its effectiveness in balancing structure preservation with appearance transfer under human perception.
Table 2. Image Appearance Transfer User Study. We show user preference rates (%) for different methods, where participants selected the result that best preserved the original scene structure while closely matching the reference style. ReStyle3D achieves the highest preference rate.
@@
Table 3. Results on Two-view Novel-view Synthesis. ReStyle3D achieves the highest scores on all metrics, indicating more accurate view synthesis and visually pleasing outputs compared to existing methods.
@@
GenWarp (Seo et al., 2024) 5122 13.503 0.465 0.435 59.965
StoryDiffusion (Zhou et al., 2024) 14.023 0.481 0.502 203.83
SDXL Inpainting (Podell et al., 2024) 16.228 0.535 0.389 89.502
ViewCrafter (Yu et al., 2024) 17.178 0.594 0.278 56.127
@@
GenWarp (Seo et al., 2024) 10242 13.491 0.565 0.440 60.540
StoryDiffusion (Zhou et al., 2024) 14.014 0.583 0.476 198.32
SDXL Inpainting (Podell et al., 2024) 16.153 0.565 0.426 89.537
ViewCrafter (Yu et al., 2024) 17.137 0.652 0.317 57.898
@@
Figure 7. Results on Video/Multi-view Appearance Transfer of ReStyle3D. We show the style images, three frames stylized by ReStyle3D, followed by a 3D reconstruction of these outputs using an off-the-shelf pipeline. Despite challenging camera motion and multiple objects in the scene, our method preserves consistent geometry and seamlessly transfers the reference style across all frames.
@@
We compare our approach to: i) SDXL inpainting model (Podell et al., 2024) with depth-conditioned ControlNet (Zhang et al., 2023c), ii) GenWarp (Seo et al., 2024), an image-based diffusion model for single view NVS, iii) StoryDiffusion (Zhou et al., 2024), a model with consistent self-attention for long-range image and video generation and iv) ViewCrafter (Yu et al., 2024), a video-diffusion model for NVS. Note that the proposed task differs from traditional NVS as it leverages geometry information from the novel view itself. We employ DUSt3R (Wang et al., 2024) to extract the correspondences and provide the initial warped image as input to all methods. As shown in Tab. 3, ReStyle3D outperforms across all metrics, achieving a superior reconstruction ability as evidenced by the best PSNR, SSIM, and LPIPS metrics. Additionally, it exhibits strong capability in extending style to unseen regions, evidenced by the lowest FID score (Fig. 10). Notably, the second best method ViewCrafter (Yu et al., 2024), requires a predefined camera trajectory as input to video diffusion and runs 10$\times$ slower than ours.
Table 4. Pose Deviation from Real-world Estimates. We measure the fraction of camera poses within certain rotation (at $5^{\circ}$, $10^{\circ}$, $15^{\circ}$) and translation (at $1\mathrm{cm}$, $2\mathrm{cm}$, $5\mathrm{cm}$) error thresholds, reporting area-under-curve (AUC) values. ReStyle3D achieves significantly higher AUC in both, showing superior multi-view geometric consistency vs. existing methods.
@@
GenWarp (Seo et al., 2024) 25.89 46.70 58.89 58.38 59.39 70.05
SDXL Inpainting (Podell et al., 2024) 34.52 52.79 66.50 61.42 65.99 74.11
ViewCrafter (Yu et al., 2024) 37.56 55.33 68.53 60.91 65.99 77.16
@@
We further evaluate the multi-view consistency of the stylized results through a proxy task Specifically, we input the original and stylized images to DUSt3R (2024) and estimate the camera poses, separately. By evaluating the agreement with the poses from the original images, we analyze whether the geometry is preserved in the stylized images. As shown in Tab. 4, our adaptive auto-regressive approach effectively mitigates inconsistencies while preserving image sharpness, significantly outperforming the baselines on all pose metrics. Figs. 10 and 11 show multi-view results, including the 3D reconstruction of stylized outputs with estimated camera poses, demonstrating both geometric and style consistency despite camera motion and multiple objects. StoryDiffusion (Zhou et al., 2024) does not support multi-view stylization.
@@
In Tab. 5(a), we run ReStyle3D without our guidance strategy and observe significant degradation in structure preservation (AbsRel from 8.34 to 16.72). In (b), removing semantic attention hurts performance on perceptual similarity w.r.t. style image, showing that both components are crucial for semantic-accurate style transfer while maintaining structural integrity. Attending unmatched instance to the style image globally provides better style fidelity compared to keeping to original image, noted as Keep. Attn..
Table 5. Ablation Study. We separately remove the guidance strategy and the semantic attention module to evaluate their impact on both structure preservation and style fidelity. Removing either significantly degrades performance, highlighting the importance of both components in achieving robust scene geometry and perceptually faithful stylization.